# ZONE: Zero-Shot Instruction-Guided Local Editing

### 28 Dec 2023 v1

@@ Abstract
Recent advances in vision-language models like Stable Diffusion have shown remarkable power in creative image synthesis and editing. However, most existing text-to-image editing methods encounter two obstacles: First, the text prompt needs to be carefully crafted to achieve good results, which is not intuitive or user-friendly. Second, they are insensitive to local edits and can irreversibly affect non-edited regions, leaving obvious editing traces. To tackle these problems, we propose a Zero-shot instructiON-guided local image Editing approach, termed ZONE. We first convert the editing intent from the user-provided instruction (e.g., “make his tie blue”) into specific image editing regions through InstructPix2Pix. We then propose a Region-IoU scheme for precise image layer extraction from an off-the-shelf segment model. We further develop an edge smoother based on FFT for seamless blending between the layer and the image.Our method allows for arbitrary manipulation of a specific region with a single instruction while preserving the rest. Extensive experiments demonstrate that our ZONE achieves remarkable local editing results and user-friendliness, outperforming state-of-the-art methods.
@@ 2.1 Generative Models for Image Manipulation
Image manipulation is a fundamental process within the realm of computer vision, involving altering images with the aid of additional conditions like textual prompts, labels, masks, or reference images. Two mainstream editing methods include Generative Adversarial Networks (GANs) and Diffusion Models (DMs). Typical image manipulation tasks comprise image-to-image translation [20, 7, 63, 26, 46, 49, 54, 10], super-resolution [31, 21, 56, 13], inpainting [41, 19, 34, 44], colorization [4, 37, 33, 53], and more. Although GAN-based methods excel when dealing with carefully curated data, they struggle with extensive and heterogeneous datasets [22, 23, 36]. To enhance generative expressiveness, [17, 50, 51, 18, 44, 58, 57] utilize DMs to achieve high-quality generation over diverse datasets. Recent research has yielded promising generation outcomes through the training or fine-tuning of large-scale text-to-image models [5, 24, 55, 35, 43, 47, 38], as well as by harnessing CLIP [42] embeddings to guide image manipulation using textual prompts [9, 25, 30]. Some prior works [1, 2, 15, 40, 52] also demonstrate the zero-shot editing capability of pretrained DMs. Similarly, our method extensively exploits a pretrained DM’s generative capability to facilitate diverse and stylized image editing. However, we uniquely explore the implicit relationship between the DM’s editing regions during generation and the whole user instructions, enabling fine-grained layer-specific positioning.
@@ 2.3 Instruction-Guided Editing
Figure 1: Overview of ZONE. (a) Three modules in ZONE. (b) Distinct difference between description-guided and instruction-guided diffusion models on cross-attention. The former usually follows a token-aware format, while the latter is edit-aware. (c) Implementation details of the modules shown in (a).
@@ 4.2 Instruction-Guided Localization
As shown in Fig. 1 (b), our key finding is that the operational mechanisms of instruction-guided and description-guided diffusion models on cross-attention exhibit a distinct difference. Specifically, we empirically demonstrate that: (i) a description-guided model displays a token-aware characteristic on its cross-attention maps, associating each input text token with a corresponding spatial structure; (ii) an instruction-guided model’s cross-attention maps share similar spatial features, demonstrating an edit-aware characteristic, being insensitive to single tokens but responsive to the overall editing intent.
@@
where the query $Q=W_{Q}\phi(z_{t})$, key $K=W_{K}c_{T}$, and value $V=W_{V}c_{T}$ are obtained with linear projections $W_{Q}$, $W_{K}$, and $W_{V}$. $M\in\mathbb{R}^{H^{{}^{\prime}}\times W^{{}^{\prime}}\times L}$ contains $L$ cross-attention maps that are correlated to the similarity between $Q$ and $K$. Typically $H^{{}^{\prime}}$ and $W^{{}^{\prime}}$ are $1/32$ of the original image size $H$ and $W$ in Stable Diffusion. For the description-guided Stable Diffusion model, let $M^{l}$ be the attention map of the $l$-th token, $l\in\{1,2,\ldots,L\}$. For the instruction-guided IP2P, $M^{l}$ shares a uniform characteristic across all tokens, concentrated directly at the edited location without token specification, as visualized in Fig. 2.
Figure 2: Token-wise cross-attention map difference. We average the cross-attention maps among all timesteps for each sample. IP2P shows consistency in all tokens, while Stable Diffusion (SD) demonstrates a one-to-one correspondence with tokens.
Based on this finding, we devise a simple yet effective localization module that semantically locates the edited region with instruction $\mathcal{T}_{I}$. Specifically, we first collect the attention maps of the denoising model of IP2P from all timesteps of the denoising process. Then, we average and resize the maps to obtain averaged attention maps $\mathcal{M}_{A}\in\mathbb{R}^{H\times W\times L}$. Note that the $L$ tokens include the “start-of-text” and “end-of-text” tokens, whose corresponding attention maps are $M^{1}$ and $M^{L}$, respectively. As depicted in Fig. 2, the attention weights (i.e., the brightness of the pixels) of these maps tend to decrease along the tokens, so we subtract the last token’s cross-attention map from the first token’s and binarize the result with a fixed threshold $T$:
@@ 4.4 Layer Blending
As shown in the second column of Fig. 3, we observe that both $\mathcal{I}_{G,d}$ and $\mathcal{I}_{L,d}$ share similar low-frequency characteristics on non-edited regions (e.g., background), but they hold different low-frequency characteristics on the edited regions (e.g., hat and the shadow below it). Therefore, we can exclude the non-edited regions and retain the edited regions by subtracting the low-frequency components between $\mathcal{I}_{L,d}$ and $\mathcal{I}_{G,d}$ in the frequency domain: $d_{s}=\mathcal{H}(f(\mathcal{I}_{L,d}))-\mathcal{H}(f(\mathcal{I}_{G,d}))$ and invert it back to the image domain to get the difference mask $\mathcal{M}_{dm}=f^{-1}(d_{s})$. The final mask $\mathcal{M}_{f}^{*}$ is then obtained by $\mathcal{M}_{f}^{*}=g(\mathcal{M}_{dm})=e(\mathcal{I}_{L,d},\mathcal{I}_{G,d})$. Finally, we get the final edited image layer $\mathcal{I}_{L}$ by $\mathcal{I}_{L}=\mathcal{I}_{sty}\odot\mathcal{M}_{f}^{*}$, and the final edited result $\mathcal{I}_{C}$ is acquired by compositing $\mathcal{I}_{G}$ and $\mathcal{I}_{L}$. The intuitive visualization of these intermediate results are shown in Fig. 3.
@@ 5.2 Comparisons
DiffEdit [8] 27.0 $\pm$ 5.6 8.8
Text2LIVE [3] 32.3 $\pm$ 4.3 17.3
Pix2Pix-Zero [40] 18.3 $\pm$ 4.3 10.4
InstructPix2Pix [5] 60.6 $\pm$ 4.9 18.9
MagicBrush [59] 50.0 $\pm$ 5.1 18.0
ZONE (Ours) 69.0 $\pm$ 3.7 26.6
@@ 6 Conclusion
In this supplementary, we first give more visualization results, then detail the datasets and the implementation, and finally state the social impact and the limitations.

### 12 Apr 2024 v2

@@ Abstract
Recent advances in vision-language models like Stable Diffusion have shown remarkable power in creative image synthesis and editing. However, most existing text-to-image editing methods encounter two obstacles: First, the text prompt needs to be carefully crafted to achieve good results, which is not intuitive or user-friendly. Second, they are insensitive to local edits and can irreversibly affect non-edited regions, leaving obvious editing traces. To tackle these problems, we propose a Zero-shot instructiON-guided local image Editing approach, termed ZONE. We first convert the editing intent from the user-provided instruction (e.g., “make his tie blue”) into specific image editing regions through InstructPix2Pix. We then propose a Region-IoU scheme for precise image layer extraction from an off-the-shelf segment model. We further develop an edge smoother based on FFT for seamless blending between the layer and the image. Our method allows for arbitrary manipulation of a specific region with a single instruction while preserving the rest. Extensive experiments demonstrate that our ZONE achieves remarkable local editing results and user-friendliness, outperforming state-of-the-art methods. Code is available at https://github.com/lsl001006/ZONE.
@@ 2.1 Generative Models for Image Manipulation
Image manipulation is a fundamental process within the realm of computer vision, involving altering images with the aid of additional conditions like textual prompts, labels, masks, or reference images. Two mainstream editing methods include Generative Adversarial Networks (GANs) and Diffusion Models (DMs). Typical image manipulation tasks comprise image-to-image translation [21, 7, 64, 27, 47, 50, 55, 10], super-resolution [32, 22, 57, 13], inpainting [42, 20, 35, 45], colorization [4, 38, 34, 54], and more. Although GAN-based methods excel when dealing with carefully curated data, they struggle with extensive and heterogeneous datasets [23, 24, 37]. To enhance generative expressiveness, [17, 51, 52, 18, 45, 59, 58] utilize DMs to achieve high-quality generation over diverse datasets. Recent research has yielded promising generation outcomes through the training or fine-tuning of large-scale text-to-image models [19, 5, 25, 56, 36, 44, 48, 39], as well as by harnessing CLIP [43] embeddings to guide image manipulation using textual prompts [9, 26, 31]. Some prior works [1, 2, 15, 41, 53] also demonstrate the zero-shot editing capability of pretrained DMs. Similarly, our method extensively exploits a pretrained DM’s generative capability to facilitate diverse and stylized image editing. However, we uniquely explore the implicit relationship between the DM’s editing regions during generation and the whole user instructions, enabling fine-grained layer-specific positioning.
@@ 2.3 Instruction-Guided Editing
Figure 1: Overview of ZONE. (a) Three modules in ZONE. (b) The distinct difference between description-guided and instruction-guided diffusion models on cross-attention. The former usually follows a token-aware format, while the latter is edit-aware. $\varnothing$ denotes the unconditional embeddings for null input. (c) Implementation details of the modules shown in (a).
@@ 4.2 Instruction-Guided Localization
As shown in Fig. 1 (b), our key finding is that the operational mechanisms of instruction-guided and description-guided diffusion models on cross-attention exhibit a distinct difference. Specifically, we empirically demonstrate that: (i) a description-guided model displays a token-aware characteristic on its cross-attention maps, associating each input text token with a corresponding spatial structure; (ii) an instruction-guided model’s cross-attention maps with unconditional embeddings share similar spatial features, demonstrating an edit-aware characteristic, being responsive to the overall editing intent.
@@
where the query $Q=W_{Q}\phi(z_{t})$, key $K=W_{K}c_{T}$, and value $V=W_{V}c_{T}$ are obtained with linear projections $W_{Q}$, $W_{K}$, and $W_{V}$. $M\in\mathbb{R}^{H^{{}^{\prime}}\times W^{{}^{\prime}}\times L}$ contains $L$ cross-attention maps that are correlated to the similarity between $Q$ and $K$. Typically $H^{{}^{\prime}}$ and $W^{{}^{\prime}}$ are $1/32$ of the original image size $H$ and $W$ in Stable Diffusion. For the description-guided Stable Diffusion model, each token corresponds to its specific attention map $M^{l}$ with text embeddings, where $l\in\{1,2,\ldots,L\}$. For the instruction-guided IP2P, we find its attention maps share a uniform characteristic with unconditional embeddings, concentrated directly at the edited location without token specification, as visualized in Fig. 2.
Figure 2: Cross-attention map difference. We average the cross-attention maps among all timesteps for each sample. IP2P shows consistency in the overall editing intent with unconditional embeddings $\varnothing$, while Stable Diffusion (SD) demonstrates a one-to-one correspondence with text tokens.
Based on this finding, we devise a simple yet effective localization module that semantically locates the edited region with instruction $\mathcal{T}_{I}$. Specifically, we first collect the attention maps of the denoising model of IP2P from all timesteps of the denoising process. Then, we average and resize the maps to obtain averaged attention maps $\mathcal{M}_{A}\in\mathbb{R}^{H\times W\times L}$. We observe that the first attention map $\mathcal{M}_{A}^{1}$ primarily emphasizes the attention weights of the non-edited region. Subsequent attention maps $\mathcal{M}_{A}^{2,...,L}$ shift attention towards the edited region. Therefore, we subtract the last cross-attention map from the first attention map and binarize each pixel $(m,n)$ with a fixed threshold $T$ to highlight the edited region and mitigate the background noise:
@@ 4.4 Layer Blending
As shown in the second column of Fig. 3, we observe that both $\mathcal{I}_{G,d}$ and $\mathcal{I}_{L,d}$ share similar low-frequency characteristics on non-edited regions (e.g., background), but they hold different low-frequency characteristics on the edited regions (e.g., hat and the shadow below it). Therefore, we can exclude the non-edited regions and retain the edited regions by subtracting the low-frequency components between $\mathcal{I}_{L,d}$ and $\mathcal{I}_{G,d}$ in the frequency domain: $d_{s}=\mathcal{H}(f(\mathcal{I}_{L,d}))-\mathcal{H}(f(\mathcal{I}_{G,d}))$ and invert it back to the image domain to get the difference mask $\mathcal{M}_{dm}=f^{-1}(d_{s})$. The final mask $\mathcal{M}_{f}^{*}$ is then obtained by $\mathcal{M}_{f}^{*}=g(\mathcal{M}_{dm})=e(\mathcal{I}_{L,d},\mathcal{I}_{G,d})$. Finally, we get the final edited image layer $\mathcal{I}_{L}$ by $\mathcal{I}_{L}=\mathcal{I}_{sty}\odot\mathcal{M}_{f}^{*}$, and the final edited result $\mathcal{I}_{C}$ is acquired by compositing $\mathcal{I}_{G}$ and $\mathcal{I}_{L}$. The intuitive visualization of these intermediate results is shown in Fig. 3.
@@ 5.2 Comparisons
DiffEdit [8] 27.1 $\pm$ 2.7 8.8
Text2LIVE [3] 33.0 $\pm$ 3.2 17.3
Pix2Pix-Zero [41] 19.2 $\pm$ 3.7 10.4
InstructPix2Pix [5] 59.8 $\pm$ 3.1 18.9
MagicBrush [60] 50.2 $\pm$ 2.9 18.0
ZONE (Ours) 69.4 $\pm$ 3.5 26.6
@@ 6 Conclusion
Acknowledgements. The work was supported by the National Key Research and Development Program of China (2023YFC3300029), Zhejiang Provincial Natural Science Foundation of China (LD24F020007), Beijing Natural Science Foundation (L223024), National Natural Science Foundation of China (62076016), “One Thousand Plan” projects in Jiangxi Province (Jxsg2023102268), Beijing Municipal Science & Technology Commission, Administrative Commission of Zhongguancun Science Park (Z231100005923035).
@@
In this supplementary, we first give more visualization results, then detail the datasets and the implementation, and finally discuss the limitations and state the social impact.
@@ C.4 Human Evaluation
C.5 Additional Ablations
We perform further ablations to demonstrate the effectiveness of each component within our proposed ZONE. First, we ablate on the $\beta$ parameter of the fused IP2P module. $\beta$ is a hyperparameter that controls the guidance strength of MagicBrush (MB) on the fused IP2P module. A higher $\beta$ emphasizes MB’s effects on the editing results (please check L302-314 in our paper for more details). We illustrate the results generated by the models under various $\beta$ in Fig. 12:
Figure 12: Ablation on $\beta$. As $\beta$ decreases from left to right, the generated result increasingly resembles that produced by InstructPix2Pix and less resembles that generated by MagicBrush. Given the limits of current instruction-guided diffusion models, we employ this character to handle different editing actions.
Then we conduct an ablation study on the components of ZONE, including SAM and the edge smoother, with human evaluation using the user preference rate (UPR) and the CLIP image similarity (CLIP-I) for metrics evaluation. As demonstrated in Tab. 5, the implementation of SAM notably boosts the user preference rate, signifying a visually discernible enhancement compared to its absence, a fact also mirrored in the CLIP-I score. Concurrently, the user preference rate exhibits a significant improvement with the integration of the Edge Smoother, underscoring its efficacy.
Component SAM Edge Smoother
w/o w/ w/o w/
CLIP-I $\uparrow$ 0.95 0.97 0.97 0.97
UPR (%) $\uparrow$ 5.4 94.6 22.7 77.3
Table 5: Ablation on ZONE’s components. We evaluate both components with human evaluation and CLIP metrics.